OpenStudy (anonymous):

how do you solve this: 2x-3 = (3+x)/2

OpenStudy (anonymous):

multiply by 2, subtract an x and add a 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

i need the process, not the answer which i already know

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

The point is, you want to get all your x terms on one side and your non-variable terms on the other

OpenStudy (anonymous):

\[2*(2x-3)=[(3+x)/2]*2\]

OpenStudy (anonymous):

then it goes to : 4x-6=3+x

OpenStudy (anonymous):

thank you so much, i understand now

OpenStudy (anonymous):

the get x by itself...so subtract -x to the left so next step: 3x=9

OpenStudy (anonymous):

do i get a medal :)

OpenStudy (anonymous):

yes! but i don't know how to give one

OpenStudy (anonymous):

i need help with this one too please: 2x-1= 1/3(5-3x) + 4

OpenStudy (anonymous):

Ok, well why don't you tell us how to start this one

OpenStudy (anonymous):

Oh, and is it \[\frac{1}{3(5-3x)}\] or \[\frac{1}{3}(5-3x)\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Ok. So which part do you want to tackle first? (There's lots of different paths to the same answer)

OpenStudy (anonymous):

Overall goal is to get all the x terms on one side and the non-x terms on the other side

OpenStudy (anonymous):

You tell me what to do, and I'll write the new equation.

OpenStudy (anonymous):

\[2x-1= \frac{1}{3}(5-3x) + 4\]

OpenStudy (anonymous):

i don't know how to start that's why i'm asking

OpenStudy (anonymous):

That's ok. Just pick a term that is in the wrong place

OpenStudy (anonymous):

We want everything with an x on the left, and everything without an x on the right

OpenStudy (anonymous):

Pick a term, any term.

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

Ah, that term is a little tricky because it's part of a factor in a product. We will need to expand that product out first before we can work with that term directly

OpenStudy (anonymous):

But that's ok!

OpenStudy (anonymous):

That just means we need to distribute the 1/3 to each of the terms in the other factor

OpenStudy (anonymous):

So what do we get when we distribute that 1/3?

OpenStudy (anonymous):

3-9x?

OpenStudy (anonymous):

no. we have to multiply each term in the left factor by 1/3 \[\frac{1}{3}(5-3x) = \frac{1}{3}*5 - \frac{1}{3}*3x\]

OpenStudy (anonymous):

Does that make sense? So what do we have when we simplify that product?

OpenStudy (anonymous):

i'm not sure...

OpenStudy (anonymous):

a(b+c) = ab + ac Right? Basic multiplicative distribution

OpenStudy (anonymous):

ok so would it be 5/3 -x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now we have: \[2x-1= \frac{5}{3}-x + 4\]

OpenStudy (anonymous):

so the whole thing would be : 2x-1 = 5/3-x+4?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Can you solve from there? or do you need more help with it?

OpenStudy (anonymous):

maybe a little more

OpenStudy (anonymous):

Ok, so pick a term that's in the wrong place

OpenStudy (anonymous):

how would i combine 5/3 + 4 that's the next step right?

OpenStudy (anonymous):

You can certainly do that yes.

OpenStudy (anonymous):

There is no 'next step'. There are a lot of ways to do it. If you want to do that next, go for it =)

OpenStudy (anonymous):

ok so how do you do it?

OpenStudy (anonymous):

How do you add \(\frac{5}{3} + 4\)

OpenStudy (anonymous):

You have to change the denominator of the 4 to match the 3 in the 5/3

OpenStudy (anonymous):

We do this using the trick of multiplying by 1.

OpenStudy (anonymous):

\(4 \times 1 = ?\)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Ok, so if we multiply by 1 we won't have changed the value, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Ok, so what is \(\frac{3}{3}\)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

So you're saying then that if I multiply 4 by 3/3 I will still have 4?

OpenStudy (anonymous):

\[4\times \frac{3}{3} = \frac{4\times 3}{3} = \frac{12}{3}\]

OpenStudy (anonymous):

And 12/3 does equal 4, but it's in a different form that is now easy to add to 5/3

OpenStudy (anonymous):

Did you follow that?

OpenStudy (anonymous):

yes so the answer would be 17/3

OpenStudy (anonymous):

Well, the sum of those two terms would be 17/3 yes

OpenStudy (anonymous):

So we have: \[2x-1= \frac{17}{3}-x\]

OpenStudy (anonymous):

then irt would be 3x-1= 17/3

OpenStudy (anonymous):

Certainly

OpenStudy (anonymous):

it would be 3x = 1 17/3?

OpenStudy (anonymous):

yeah, but I dun like mixed numbers

OpenStudy (anonymous):

Can you do the same trick to the 1 we did to the 4 to make it easy to add to the 17/3

OpenStudy (anonymous):

isn't it 20/9?

OpenStudy (anonymous):

x = 20/9 is correct yes.

OpenStudy (anonymous):

you don't divide that by 3?

OpenStudy (anonymous):

Huh? we did already

OpenStudy (anonymous):

It was 3x = 20/3

OpenStudy (anonymous):

oh ok i thought it was 3x = 1 17/3? and then you divide by 3 on both sides to get x

OpenStudy (anonymous):

\(3x = 1+\frac{17}{3}\)\[\implies 3x = 1\times \frac{3}{3} + \frac{17}{3}\]\[\implies 3x = \frac{1\times 3}{3} + \frac{17}{3}\]\[\implies 3x = \frac{3}{3} + \frac{17}{3}\]\[\implies 3x = \frac{20}{3}\]

OpenStudy (anonymous):

Then you divide by 3 on both sides

OpenStudy (anonymous):

Follow all that?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

but how do you divide 3 on both sides, that what i want to know

OpenStudy (anonymous):

dividing by 3 is the same as multiplying by 1/3

OpenStudy (anonymous):

wouldn't the answer be different?

OpenStudy (anonymous):

nope. \[3x \times \frac{1}{3} = \frac{20}{3} \times \frac{1}{3}\] \[x = \frac{20}{9} \] \[3x \div 3 = \frac{20}{3} \div 3\] \[x = 6.\bar{6}\bar{6} \div 3 = 2.\bar{2}\bar{2} = \frac{20}{9}\]